# Unitary quantum gravitational physics and the CMB parity asymmetry

### v1

@@ Appendix C Appendix C: DSI angular power spectrum
$\Delta\mathcal{P}_{v}=\left(1-n_{s}\right)\operatorname{Re}\left[\frac{1}{H_{3/2}^{(1)}\left(\frac{k}{k_{\ast}}\right)}\frac{\partial H^{(1)}_{\nu_{s}}\left(\frac{k}{k_{\ast}}\right)}{\partial\nu_{s}}\Bigg{|}_{\nu_{s}=\frac{3}{2}}\right]$ (22)
Here, $H^{(1)}_{\nu_{s}}\left(z\right),j_{\ell}\left(z\right)$ are the Hankel and Bessel functions of the first kind, $k_{s}=7\times 10^{-5}\,{\rm Mpc}^{-1}$ and $k_{c}=0.02k_{\ast}$ is the cut-scale we impose as (22) is accurate enough for $-\ell$ or large angular scales. This cut-off scale is also related to the coarse gaining scale of stochastic inflation Gaztañaga and Sravan Kumar (2024) which determines the large wavelength modes that have already become classical on the onset of inflation while the mode $k_{\ast}$ exiting the horizon. In DSI, these large wavelength modes create parity asymmetric inhomogeneities that correct the FLRW spacetime on large scales. The parity asymmetry ceases to exist towards the small scales because the $\mathcal{P}\mathcal{T}$ symmetry recovered towards the short distance scales. This can be seen in Fig. 11 of Gaztañaga and Sravan Kumar (2024). An intuitive explanation for this is that large wavelength modes experience quantum gravitational effects more than those of short wavelength modes.

### v2

@@ Appendix C DSI angular power spectrum
$\Delta\mathcal{P}_{v}=\left(1-n_{s}\right)\operatorname{Re}\left[\frac{2}{H_{3/2}^{(1)}\left(\frac{k}{k_{\ast}}\right)}\frac{\partial H^{(1)}_{\nu_{s}}\left(\frac{k}{k_{\ast}}\right)}{\partial\nu_{s}}\Bigg{|}_{\nu_{s}=\frac{3}{2}}\right]$ (28)
Here, $H^{(1)}_{\nu_{s}}\left(z\right),j_{\ell}\left(z\right)$ are the Hankel and Bessel functions of the first kind, $k_{s}=7\times 10^{-5}\,{\rm Mpc}^{-1}$ and $k_{c}=0.02k_{\ast}$ is the cut-scale we impose as Eq.28 is accurate enough for $-\ell$ or large angular scales. This cutoff scale is also related to the coarse-graining scale of stochastic inflation (Gaztañaga and Kumar (2024)), which determines the large wavelength modes that have already become classical on the onset of inflation while the mode $k_{\ast}$ exits the horizon. In DSI, these large-wavelength modes create parity-asymmetric inhomogeneities that correct the FLRW spacetime on large scales. The parity asymmetry ceases to exist towards the small scales because the $\mathcal{P}\mathcal{T}$ symmetry is recovered towards the short-distance scales. This can be seen in Fig. 11 of Gaztañaga and Kumar (2024). An intuitive explanation for this is that large wavelength modes experience quantum gravitational effects more than those of short wavelength modes. For the small scales $k>k_{c}$, the predictions of DSI become almost the same as with the (near) scale-invariant power spectrum as contributions from $\Delta\mathcal{P}_{v}$ term become negligible.